OpenStudy (anonymous):

Graph the function D=V1.5h which gives the approximate distance d in miles to the horizon, where h is the height of the viewer’s eyes above the ground in feet. Approximately how high above the ground must a person be if the horizon is 17.2 miles away?The v is actually square root

OpenStudy (phi):

\[ D= \sqrt{1.5h} \] to graph this, pick a number for h, for example h=0 put that number into the formula \[ D= \sqrt{1.5\cdot 0} = \sqrt{0}= 0\] (use a calculator, as most numbers are not as easy to do as 0) you now have a pair of numbers: h=0, D=0 let the h be the x-axis, and D be the y-axis. plot the point (0,0) now do h=1 \[ D= \sqrt{1.5\cdot1}= \sqrt{1.5}= 1.22\] plot the point (1,1.22) do that for more points and connect the dots.